OpenStudy (anonymous):

1. the melting point of sodium is 98 degrees celsius and the melting point of zinc is 420 celsius. How much hotter is the melting point of zinc than that of sodium? 2. The average temperature of Venus is 480 C. the average temperature on pluto is -230c. How much warmer is Venus than pluto? 3. Aluminum melts at 659c and copper melts at 1083c what is the difference between the melting points of aluminum and copper? 4. The boiling point of water is 212 degrees far. Propane boils at about -44 f. how much hotter is the boiling point of water than of propane

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

??

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

You're doing arithmetic here. All the units are the same, so you don't have to convert, just add and subtract as necessary. Are you having trouble setting up the problems so you can solve them?

OpenStudy (anonymous):

1) 420 - 98 = 322 2) 480 - (-230) = 480 + 230 = 710 3) 1083 - 659 = 424 4) 212 - (-44) = 212 + 44 = 256

OpenStudy (anonymous):

As petiteJiji ands stellarpiper stated. These are simple arithmetic problems which do not involve any conversions.